OpenStudy (anonymous):

Solve the equation. Check for extraneous solutions. 2|7-7x|=2x+4

OpenStudy (anonymous):

Divide both sides by 2 first.

OpenStudy (anonymous):

Show me what you come up with after that step I just told you.

OpenStudy (anonymous):

7-7x=x+4?

OpenStudy (anonymous):

why is the +4 still there when you divided both sides by 2.

OpenStudy (anonymous):

?*

OpenStudy (anonymous):

no, dont forget to divide everything!

OpenStudy (anonymous):

so that side would be x+2 then

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

Now when you get rid of the absolute value signs, you get two solution sets.

OpenStudy (anonymous):

\[7-7x=x+2\] or \[7-7x=-(x+2)\]

OpenStudy (anonymous):

Do you understand that or have you learnt that in class?

OpenStudy (anonymous):

yeah I get that

OpenStudy (anonymous):

Okay solve for x as normal for the first equation first.

OpenStudy (anonymous):

Tell me what you get for x in the first equation.

OpenStudy (anonymous):

wait wouldn`t the x`s cancel each other out? I`m so lost

OpenStudy (anonymous):

You don't know how to solve this simple equation? \[7-7x=x+2\]

OpenStudy (anonymous):

No need to be mean we are all at diffrent levels of math here/

OpenStudy (anonymous):

I'm asking her so then if she says yes, I'm going show her step by step.

OpenStudy (anonymous):

No I don`t know how

OpenStudy (anonymous):

Okay. Then Imagine x is the number 1.

OpenStudy (anonymous):

and 7x is the number 7

OpenStudy (anonymous):

Don't look at the pronumeral. Just look at the number beside it.

OpenStudy (anonymous):

here is what you do from here on out. minus 2 from both sides. 5-7x=x Then add 7x to both sides. -> 5=8x divide by 8 both sides -> 5/8=x

OpenStudy (anonymous):

@AIK Good job. I'm showing her step by step so she can understand.

OpenStudy (anonymous):

So it would be 5/8

OpenStudy (anonymous):

Do you know how to get that? Because if you just say "Yes, I know how to get it", then it's your fault if you get a really low mark in maths.

OpenStudy (anonymous):

@Brittanyy! Do you know what "collecting like terms" is?

OpenStudy (anonymous):

Yes I really do get it. You minus 2 from both sides to cancel the 2 on the right side and get x by itself then add 7x to both sides to get x on one side then divide by 8 to find x. Thanks so much!

OpenStudy (anonymous):

Then do the second equation by yourself and tell me the answer for the second one.

OpenStudy (anonymous):

If you understand this, then I'm sure you can finish off the second equation and tell me the answer.

OpenStudy (anonymous):

@Brittanyy! Do you have the answer now?

OpenStudy (anonymous):

Remember everything in a equation is related. So x+2=2 -> you can not only minus 2 from one side, you have to from both x+2-2=2-2 -> x=0 It is the same for division and multiplication ->like terms is gathering all numbers of a same kind together. It is to make the equation easier by getting a number=variable P.S when doing l7-7xl do not forget to do both of the possible equations and check the answers

OpenStudy (anonymous):

Ah god, I have a feeling that you're giving answers and then letting the receiver fail in a maths exam later on. But who am I to judge because I'm not going to be the one who's going to fail.